# Logical computation demonstrated with a neutral atom quantum processor

### v2

@@ IV Encoded $24$-qubit cat state
Figure 5: A portion of the encoded circuit from Fig. 4 prepares an encoded four-qubit cat state. The gates in this circuit need to be scheduled. Two-qubit CZ gates are applied in the interaction zone, up to eight gates in parallel, and all the one-qubit gates on a particular qubit are applied at its initial location in the storage zone. One-qubit gates on a subset of qubits in the same row can be applied in parallel. Two atoms can be moved in parallel, but their paths cannot cross. Note that in this example, for clarity the two code blocks are arranged adjacent to each other in the storage zone, whereas in Fig. 4 they are interleaved.
Figure 6: Circuit to prepare a $24$-qubit cat state, unencoded. One-qubit gates are applied at an atom’s home register location.
The cat state has genuine multipartite entanglement, meaning that it is entangled across every bipartition of the $n$ qubits. Any state that has fidelity greater than $1/2$ with the ideal cat state also has genuine multipartite entanglement GS (10). However, states with $\leq 1/2$ fidelity may not even be entangled; the classical mixture $\tfrac{1}{2}({{|0^{n}\rangle}\!{\langle 0^{n}|}}+{{|1^{n}\rangle}\!{\langle 1^{n}|}})$ has fidelity $1/2$ with the ideal cat state, but no entanglement. Our experiments, as well as HDHL (24); RAC+ (24), measure $X$ basis and $Z$ basis error rates, $p_{X}$ and $p_{Z}$. The fidelity to the ideal cat state is between $1-p_{X}-p_{Z}$ and $1-\max(p_{X},p_{Z})$, essentially because possible $Y$ errors contribute to both $p_{X}$ and $p_{Z}$; this is the reason for the fidelity ranges in Table 1.
We find that for the encoded cat state, using error detection and loss correction, $p_{X}+p_{Z}=26.6^{+3.3}_{-3.1}\%$, while for the unencoded baseline $p_{X}+p_{Z}=42.0^{+11.3}_{-10.1}\%$. The results are summarized in Fig. 1 and Table 3. While the encoding is clearly beneficial, interestingly, it helps reduce $p_{X}$ much more than $p_{Z}$. The baseline error bar is large because rather than averaging over all trials, we chose for a challenging baseline the best of $27$ separate runs, each consisting of 300300300300400400400400 unencoded trials with $X$ and $Z$ basis measurement; Fig. 7 gives more details.
Figure 8 shows how conditioning on at most $k$ lost qubits, for various values of $k$, allows a smoother interpolation, in both acceptance probability and total error probability, between the different results in Table 3. Finally, Fig. 9 shows the different machine runs that combine to give the data in Table 3, plotted conditioned on no initial atom loss. The spread and correlations in the figure suggest some machine parameter drift.
Figure 7: Each trial of the unencoded baseline $24$-qubit cat state experiment included two cat states, one measured in the $X$ basis and the other measured in the $Z$ basis. Shown here are the baseline runs, each consisting of 300300300300400400400400 trials. Only cat states with no qubit loss are included, on average $135.3$ and $161.0$ trials for $X$ and $Z$ basis measurements, respectively. The high variance in the observed error rates complicates reporting a single baseline $X$ plus $Z$ basis total error rate. To be conservative, disadvantaging the encoded cases, we decided to report the best single run as a baseline. The red data point has the least total error. In the $X$ basis, $26$ out of $100$ cat states had a logical error, and in the $Z$ basis, $19$ out of $121$ cat states had a logical error. The baseline error bars in Fig. 1 are large because they are based on only this run, rather than on the average of all the runs in this plot.
Figure 8: Data from the encoded $24$-qubit cat state experiment, selecting only trials within specified loss criteria. In general, attempting to correct for more loss leads to a higher acceptance probability, but also a higher error rate. In each trial, an atom may be missing before the circuit even begins, a loading loss, or it may be lost during the circuit. The red points show the total $X$ basis plus $Z$ basis error rate, plotted against the average of the $X$ and $Z$ acceptance probabilities, based on either all 29 7602976029\,76029 760 trials, only those 17 0081700817\,00817 008 trials with no initial loss, or only those $858$ trials with no loss at all. For the blue points, we select only trials in which the total number of lost qubits, including load loss and circuit loss, is at most a certain threshold. Note that Fig. 9 only includes trials with no initial loss, and the acceptance probabilities in Fig. 9 are conditioned on no initial loss, so the normalization is different. (Figure 7 for the baseline only includes trials with no loss at all.)
Table 3: Data for the $24$-qubit cat state experiments. An experiment is accepted if it is neither prerejected nor postrejected. For the encoded experiments, a trial is “prerejected” if it is rejected during the initial state preparation because one of the ${|\overline{00}\rangle}=\tfrac{1}{\sqrt{2}}({|0^{4}\rangle}+{|1^{4}\rangle})$ states fails verification. A surviving trial is “postrejected” if a code block cannot be decoded, either because of a detected error or because of two or more lost qubits in the block. We have also filtered out the encoded data with no initial losses, and the data with no losses at all. For the unencoded baseline, a trial is prerejected if there are any initial losses and a surviving trial is postrejected if there are any circuit losses. As explained in Fig. 7, we have selected the best of the unencoded runs, in order to create a challenging baseline.
  Experiment Basis Trials Prerejected Postrejected $\Pr[\text{accept}]$ Errors Error rate
Cat-$24$ encoded: all data $X$ 14 4001440014\,40014 400 10 1591015910\,15910 159 $3026$ $8.4^{+0.5}_{-0.4}\%$ 148 $12(2)\%$
$Z$ 15 3601536015\,36015 360 11 0901109011\,09011 090 3283 $6.4(4)\%$ 170 $17(2)\%$
no initial loss $X$ $8433$ 5371 2122 $11.1(7)\%$ 115 $12(2)\%$
$Z$ $8575$ 5589 2185 $9.3(6)\%$ 115 $14^{+3}_{-2}\%$
no loss $X$ 395 66 221 $27^{+5}_{-4}\%$ 9 $8^{+6}_{-4}\%$
$Z$ 463 98 220 $31(4)\%$ 2 $1^{+3}_{-1}\%$
Cat-$24$ unencoded baseline $X$ 300 87 113 33(5)26 $26^{+8}_{-9}\%$
$Z$ 300 82 97 $40^{+6}_{-5}\%$ 19 $16^{+7}_{-6}\%$
(a) $X$ basis runs
(b) $Z$ basis runs
Figure 9: Encoded $24$-qubit cat state data from 45 $X$ basis measurement runs, and 48 $Z$ basis measurement runs. This data was taken overnight, while the machine ran unattended. Each blue point gives the acceptance probability, conditioned on no initial loss, and logical error rate for a run of, on average, $187$ trials for the $X$ basis data, and $179$ trials for the $Z$ basis data. (We attempted $320$ trials, but some trials failed to load all atoms.) For clarity, only some error bars are shown. The black squares are the averages over all the data, the numbers we report. We interpret the negative correlation between acceptance probability and logical error rate as evidence of some machine parameter drift between runs, despite automatic recalibrations.
V Encoded Bernstein-Vazirani algorithm
Consider a black box oracle $\{0,1\}^{n}\rightarrow\{0,1\}$, with a secret string $s\in\{0,1\}^{n}$, that on query $x$ returns $s\cdot x=s_{1}x_{1}\oplus\cdots\oplus s_{n}x_{n}$, with $\oplus$ denoting addition modulo $2$. Classically, $n$ queries to the oracle are required and sufficient to learn $s$. The Bernstein-Vazirani algorithm BV (97); CEMM (98) makes a single query to the quantum version of this oracle, $U_{s}=\sum_{x\in\{0,1\}^{n},b\in\{0,1\}}{{|b\oplus(s\cdot x)\rangle}\!{\langle b|}}\otimes{{|x\rangle}\!{\langle x|}}$, to learn $s$ through the following steps:
${|-,+^{n}\rangle}$ $\overset{U_{s}}{\longmapsto}\tfrac{1}{\sqrt{2^{n}}}\sum_{x\in\{0,1\}^{n}}{(-1)^{s\cdot x}{|-,x\rangle}}$ $\overset{I\otimes H^{\otimes n}}{\longmapsto}{|-,s\rangle}\enspace.$
This classic algorithm was designed to prove an oracle separation between the classical randomized and quantum polynomial-time complexity classes $\mathsf{BPP}$ and $\mathsf{BQP}$.
[…]
@@ V Encoded Bernstein-Vazirani algorithm
Figure 10: The $n=19$ Bernstein-Vazirani algorithm involves $n+1=20$ logical qubits. In our experiment, we test the most challenging case, in which the secret string is $1^{n}$, so there are $n$ logical CNOT gates. Encoded into the $\llbracket 4,1,2\rrbracket$ code, each logical qubit becomes a block of four atoms. There are $4(n+1)$ atoms and $26n+$ 2 physical CZ gates. Four code blocks are arranged in each of the storage zone’s top five rows. The gates are scheduled so that the first block, once initialized, stays in the interaction zone. As transversal CZ gates are applied into block $j$, block $j+1$ is initialized.
@@
Figure 2 summarizes the results, with more details in Table 4. The encoded algorithms have a statistically significantly higher success probability than the unencoded baseline, albeit with a lower acceptance rate.
Table 4: Data for the Bernstein-Vazirani experiments. The encoded experiment has no preselection. A surviving trial is rejected if one of the last $n$ code blocks cannot be decoded, either because of a detected error or from too many lost qubits. (The first code block is ignored.) For the unencoded baselines, a trial is rejected if a measured qubit is lost.
      All data
$n$ Version Trials Rejected $\Pr[\text{accept}]$ Errors Error rate
7 encoded 1200 409 $66(3)\%$ 18 $2.3^{+1.2}_{-0.9}\%$
baseline 600 224 $63(4)\%$ 57 $15^{+4}_{-3}\%$
11 encoded 1200 547 $54(3)\%$ 22 $3^{+2}_{-1}\%$
baseline 600 196 $67(4)\%$ 65 $16^{+4}_{-3}\%$
15 encoded 1200 740 $38(3)\%$ 40 $9^{+3}_{-2}\%$
baseline 600 243 $59(4)\%$ 73 $20(4)\%$
19 encoded 1800 1359 $25(2)\%$ 43 $11(3)\%$
baseline 800 457 $43(3)\%$ 90 $26^{+5}_{-4}\%$
23 encoded 2400 1993 $17^{+2}_{-1}\%$ 67 $16^{+4}_{-3}\%$
baseline 3160 2379 $25^{+2}_{-1}\%$ 212 $27(3)\%$
27 encoded 26 0402604026\,04026 040 23 7892378923\,78923 789 $8.6(3)\%$ 498 $(2)\%2$
baseline 15 6001560015\,60015 600 12 8771287712\,87712 877 $17.5(6)\%$ 857 $31(2)\%$
Trials with no initial loss
$n$ Version Trials Rejected $\Pr[\text{accept}]$ Errors Error rate
7 encoded 732 233 $68(3)\%$ 7 $1.4_{-0.8}^{+1.3}\%$
baseline 425 76 $82_{-4}^{+3}\%$ 30 $9(3)\%$
11 encoded 471 181 $62(4)\%$ 4 $1.4_{-1.0}^{+1.8}\%$
baseline 540 142 $74(4)\%$ 59 $15_{-3}^{+4}\%$
15 encoded 543 308 $43(4)\%$ 15 $6_{-3}^{+4}\%$
baseline 526 176 $67(4)\%$ 66 $19(4)\%$
19 encoded 537 369 $31(4)\%$ 16 $10_{-4}^{+5}\%$
baseline 623 289 $54(4)\%$ 81 $24_{-4}^{+5}\%$
23 encoded 800 647 $19(3)\%$ 25 $16_{-5}^{+6}\%$
baseline 2173 1409 $35(2)\%$ 195 $26(3)\%$
27 encoded 4981 4443 $10.8_{-0.8}^{+0.9}\%$ 104 $19(3)\%$
baseline 10 3581035810\,35810 358 7695 $25.7(8)\%$ 797 $30(2)\%$
@@
Computing the success probability conditioned on acceptance makes sense if we treat the Bernstein-Vazirani algorithm as a generic quantum algorithm, as done above. However, it was originally proposed as a quantum query algorithm, in which minimizing the number of oracle queries was of key importance. In this light, we can look at the data and ask whether with only one oracle query allowed (one trial), the encoded circuit still beats the unencoded circuit. In short, the answer is probably not. In this case, neither the encoded nor unencoded algorithms can condition on acceptance, and with only one trial it requires guessing the secret string. In particular, with a lost unencoded qubit or an undecodable code block the probability of guessing the corresponding bit is only $1/2$, and the expected Hamming distance of the guess away from the secret string increases by $1/2$. As shown in Table 5, if you are only allowed to make one query to either the unencoded or encoded oracles, and if your goal is to maximize the probability of guessing the secret string $s$, then for $n\geq 11$ it is better to use the unencoded oracle. On the other hand, if your goal is to minimize the expected Hamming distance of your guess from $s$, then it is better in all cases to use the encoded oracle.
These statements are not contradictory. They hold because the failure modes of the encoded and unencoded algorithms are quite different, as shown in Fig. 11. Intuitively, an error on the ancilla qubit in the unencoded algorithm typically results in a string like $1^{n/2}0^{n/2}$ and contributes on average $n/2$ to the expected Hamming distance from $s$. On the other hand, an error on the ancilla code block in the encoded algorithm often results in a string like $1^{n/2}\star^{n/2}$, where $\star$ denotes an undecodeable code block where one must guess a bit of $s$; this contributes only an average of $n/4$ to the expected Hamming distance. Typical errors in the encoded algorithm are heralded, so they contribute less to the expected Hamming distance error metric.
@@
One query to either quantum oracle is much better than one query to the classical oracle. One classical query can only return one bit of $s$, so the probability of guessing $s$ is $\Pr[\text{guess}]=1/2^{n-1}$ and the expected Hamming distance to $s$ is $\operatorname{\operatorname{E}}[\text{Hamming}]=\tfrac{n-1}{2}$.
@@
Table 5: Performance of unencoded and encoded Bernstein-Vazirani algorithms with a single oracle query. $\Pr[\text{guess}]$ is the probability of guessing the secret string based on the measured outputs, and $\operatorname{\operatorname{E}}[\text{Hamming}]$ the expected distance of the guess away from the secret string. This analysis is based on all trials, including ones with initial loss. For a classical oracle, $\Pr[\text{guess}]=1/2^{n-1}$ and $\operatorname{\operatorname{E}}[\text{Hamming}]=\tfrac{n-1}{2}$.

### v3

@@ Abstract
Figure 8: The Bernstein-Vazirani algorithm for a secret string $s\in\{0,1\}^{n}$ uses $n+1$ logical qubits. a, In the most challenging case, $s=1^{n}$ and there are $n$ logical CNOT gates. b, Encoded into the $\llbracket 4,1,2\rrbracket$ code, there are $4(n+1)$ atoms and $26n+$ 2 CZ gates. c, The gates are scheduled so that the first block, once initialized, stays in the interaction zone. As transversal CZ gates are applied into block $j$, block $j+1$ is initialized.
@@
Figure 4 summarizes the results. The encoded algorithms have at least a 1.4×1.4$\times$1.4 × lower error rate than the unencoded baseline, albeit with a lower acceptance rate.
@@
Computing the success probability conditioned on acceptance makes sense if we treat the BV algorithm as a generic quantum algorithm. However, it was originally proposed as a quantum query algorithm, in which minimizing the number of oracle queries was key. In this light, we can ask whether with only one query (trial), the encoded circuit beats the unencoded one. When the algorithms cannot condition on acceptance, it requires guessing $s$. With a lost unencoded qubit or an undecodable code block the probability of guessing the corresponding bit of $s$ is $1/2$, and the expected Hamming distance from $s$ increases by $1/2$. As shown in Table 3, if the goal is to maximize the probability of guessing $s$, based on one query to either the unencoded or encoded oracles, then for $n\geq 11$ the unencoded oracle is better. On the other hand, if the goal is to minimize the expected Hamming distance to $s$, then it is better to use the encoded oracle. These statements can both hold because the failure modes of the encoded and unencoded algorithms are different (Fig. 9). Intuitively, an error on the ancilla qubit in the unencoded algorithm typically results in a string like $1^{n/2}0^{n/2}$, contributing $n/2$ to the distance from $s$. However, with the encoded algorithm a faulty ancilla block more likely leads to $1^{n/2}\star^{n/2}$, where $\star$ denotes an undecodeable block; this contributes only $n/4$ to the expected distance. Typical errors in the encoded algorithm are heralded, so they contribute less to the expected distance.
@@
Either quantum oracle is much better than the classical oracle. One classical query can only return one bit of $s$, so the probability of guessing $s$ is $\Pr[\text{guess}]=1/2^{n-1}$ and the expected distance to $s$ is $\operatorname{\operatorname{E}}[\text{Hamming}]=\tfrac{n-1}{2}$.
@@
Table 3: Performance of unencoded and encoded Bernstein-Vazirani algorithms with one oracle query. $\Pr[\text{guess}]$ is the probability of guessing the secret $s$, and $\operatorname{\operatorname{E}}[\text{Hamming}]$ the expected distance from $s$. For a classical oracle, $\Pr[\text{guess}]=1/2^{n-1}$ and $\operatorname{\operatorname{E}}[\text{Hamming}]=\tfrac{n-1}{2}$.
@@
Outlook
In this work, we have created large entangled logical states and implemented an algorithm fault tolerantly on encoded qubits, with better-than-physical performance. These results highlight the features and promise of neutral atoms as a platform for reliable fault-tolerant quantum computing. They are enabled by the combination of scalability and high-fidelity operations inherent to neutral atoms, with full qubit connectivity.
All of the experiments highlight the usefulness of loss conversion and imaging. The encoded experiments demonstrate that logical circuits can correct for and operate effectively with atom loss, which will be a key feature of large-scale neutral atom quantum computers.
This work, and other recent work with neutral-atom qubits BEG+ (24); RRJ+ (24), superconducting qubits GC (24) and trapped ion qubits PSR+ (24); RAC+ (24), highlights a critical advance in the field of quantum computation: the transition from noisy, intermediate-scale quantum computations with physical qubits to computational advances achieved with logical qubits. After a relatively short time from its inception, the neutral atom quantum computing platform has shown significant advancement. We expect further advances in 2Q gate fidelities and significant scale up towards 10 0001000010\,00010 000 qubits, including mid-circuit measurement NC+ (23) and continuous atom reloading NKC+ (24). Larger scales will unlock nonlocal error-correcting codes with high encoding rates XZZ+ (24). The combination of a hardware-optimized qubit virtualization system, complete with efficient error correction, with a programmable neutral atom quantum processor is a promising path to large-depth, logical computation for scientific quantum advantage.
Acknowledgments
Appendix A Methods
A.1 Register loading and atom reuse
We load atoms into the register using the methods described in NKC+ (24), providing more than enough qubits for the circuit (typically filling 320 register sites to approximately 99% filling fraction over up to 14 loading cycles). The register sites are configured in eight pairs of columns, matching the shape of the IZ. This provides “highways” through which atoms can be moved with less risk of disturbing other atoms in the register. By loading more register sites than are used in a given circuit, we can improve data rates by reusing atoms in multiple circuit instances. To do this, we determine which atoms were lost in a circuit, and refill those sites using extra atoms, then perform the rest of our state preparation protocol (described in the following section). This can be repeated as long as atoms remain in the register.
A.2 State preparation and measurement
Our platform benefits from the ability to determine successful loading of our desired qubit array before circuit execution, and to separately determine the survival and final state of an atom after the circuit. We achieve this by performing three images per circuit instance, in addition to images used to initially arrange atoms into the qubit register. A first “preselect” image determines whether the atoms were successfully loaded into the register before executing the circuit. A second “readout image” comprises the state readout of the circuit. A third “survival” image determines whether an atom is still present after the circuit. In all images, we measure the presence of an atom in state ${|1\rangle}$. The pre-select and survival images are combined with optical pumping into ${|1\rangle}$, so only determine the site occupancy. In the readout image, the optical pumping is omitted and state selectivity is maintained. Following the preselect image, we perform gray-molasses cooling and optically pump into state ${|0\rangle}$.
Aspects of our imaging procedure have been described in NC+ (23); NKC+ (24). The imaging is performed in a cavity-enhanced 784 nm wavelength 3D lattice by collecting the photons scattered from the cycling transition from ${|1\rangle}$ to ${}^{3}P_{1}{|3/2,3/2\rangle}$. Optical pumping and gray-molasses cooling are also performed in the cavity-enhanced lattice. Between imaging and circuit execution, atoms are transferred between the lattice and register tweezers.
Typical classification errors, loss probabilities, and spin-flip probabilities for the imaging are provided in NKC+ (24). Here, we also characterize state preparation and measurement errors in a manner relevant to our circuit execution by simply preparing either state ${|0\rangle}$ or ${|1\rangle}$, and then measuring the state of the atoms, conditioned on the presence of an atom in the survival image. We obtain combined preparation and readout errors for ${|0\rangle}$ and ${|1\rangle}$ of $1.1(1)\times 10^{-3}$ and $1.6(1)\times 10^{-3}$, respectively.
Histograms of the recorded photon counts provide insight into the nature of the errors (Fig. 10). For both prepared states, we see clear evidence of a two-part distribution, indicating that the dominant contribution to the error is that the atom actually ends up in the wrong state, rather than overlap of the photon count distribution of dark and bright atoms. This is likely due to a combination of imperfect optical pumping and spin flips that occur within the image when measuring ${|1\rangle}$. (Spin flips are primarily due to population of ${}^{3}P_{1}$, so are suppressed when measuring the dark state ${|0\rangle}$.)
Figure 10: a, Timing diagram of imaging and optical pumping used for circuit execution. Optical pumping is used both to provide state insensitivity for preselect and survival images, and for state preparation. b, Photon count histograms recorded after preparing ${|0\rangle}$ (left) and ${|1\rangle}$ (right). The dashed vertical line indicates the count threshold used to distinguish the two states. Error rates for ${|0\rangle}$ and ${|1\rangle}$ are 0.0011(1) and 0.0016(1), respectively.
A.3 Atom movement
Arbitrary gate connectivity in our system is obtained by moving selected atoms from the register into an interaction zone (IZ) where two-qubit gates are performed (see Fig. 1). Movement is performed using a set of tweezers generated by a pair of crossed AODs. Atoms within a single row can be moved in parallel, provided that the move profile does not involve parasitic tweezer spots generated by intermodulation tones crossing the position of the intentional tweezers used for movement. In practice, we guarantee this condition by requiring parallel moves of more than two atoms to be shape-preserving (stretching and translation of the tweezer locations only).
Moves are broken up into straight segments, with the movement velocity along each segment following a sinusoidal profile, with a total duration scaled to maintain a fixed peak acceleration (as the square root of the length of the segment). Move durations are $0.12\sqrt{L}$ ms, where $L$ is the length of the move in units of the 3.3 $\mu$m grid spacing of the register. Atoms are handed between static and rearrangement tweezers by ramping the depth of the rearrangement tweezers over 0.4 ms. For up to 50 moves, we observe a typical loss probability of $5(1)\times 10^{-4}$. For more moves, loss per move increases. By performing sideband spectroscopy on the 1S0 to 3P0 transition after a variable number of moves, we find a heating rate of 0.02(5) quanta along the propagation direction of our clock addressing laser. This heating rate is too low to explain the increased rate of atom loss above 50 moves, making axial heating a more likely culprit for the excess loss.
A.4 Single-qubit gate performance
Our single-qubit (1Q) gate scheme is described in detail in Ref. MSS+ (24), and in that work characterized within a small array. These gates are performed using two-photon Raman transitions driven by two independently controlled and site-resolved laser beams detuned roughly 10 GHz from the 3P1, $F=1/2$ manifold. Single-qubit operations are performed in parallel on rows of atoms within the register, with the pulse area, phase, and presence of the gates controlled individually on each site. We calibrate an $S_{X}$ gate on each site of the register, and perform virtual $Z$ rotations of arbitrary angle by changing the relative phase of the two drive lasers on subsequent gates.
Figure 11: Characterization of single-qubit gates. a, Array-averaged results of Clifford randomized benchmarking, indicating a per-Clifford-gate error of $6.7\times 10^{-4}$. b, Characterization of the spread in per-Clifford-gate errors on different sites within six rows sampled within the register.
Here, we characterize the 1Q gates in the larger register array used in this work. As in MSS+ (24), we perform Clifford randomized benchmarking. We consider qubits occupying six rows that span those used in our circuits and extract an average error rate per Clifford gate of $6.7\times 10^{-4}$ (see Fig. 11a). The spread in gate fidelity for different qubit sites is illustrated in Fig. 11b.
In addition to errors occurring during gates, idle errors can also be relevant. Using Ramsey spectroscopy, we measure a root-mean-squared spread in the qubit frequencies of 0.52 Hz after software compensation of shifts due to magnetic field gradients and average light shifts. The ensemble-averaged coherence exhibits a Gaussian $1/e$ decay time of 3.5(17) seconds. Uncorrected shifts are likely dominated by inhomogeneous light shifts from the trapping tweezers, which we measure to be typically on the Hz scale for the register tweezers. Additionally, the rearrangement tweezers lead to shifts on the several-Hz scale, which we calibrate and cancel in software. These shifts are not fully understood, but may result from finite helicity in the tweezer polarization. These frequency shifts are small enough that we do not apply dynamical decoupling sequences for the circuits in this work.
A.5 Two-qubit gate implementation and performance
Two-qubit controlled-phase (CZ) gates are performed in the IZ on up to eight pairs of atoms in parallel, using the same scheme described in MSS+ (24). These gates are mediated by excitation to a highly-excited Rydberg state ($65^{3}\mathrm{S}_{1}$). The sites of the IZ are spaced such that atoms occupying the same pair of sites are in the so-called “Rydberg blockade” limit, where the strength of interactions between the atoms exceeds other relevant energy scales, while atoms in different pairs interact weakly enough to be treated independently. All operations in the IZ are applied in parallel to all atoms within the IZ.
Atoms are excited to the Rydberg state using a sequential two-photon transition. State-selectivity is provided by the first stage of the two-photon transition, from 1S0, $m_{F}=-1/2$ to the metastable 3P0, $m_{F}=1/2$ state. Polarization selectivity, along with the narrow linewidth of the transition and slow Rabi rate (10 kHz) relative to the qubit frequency (376 kHz) ensure that only one qubit state is excited to the metastable state. From the metastable state, atoms are excited to the Rydberg state using a UV laser pulse (total duration 110 ns) that implements a gate similar to the time-optimal controlled-phase (CZ) gate described in JP (22); PWJ+ (22); EBK+ (23). While the IZ traps are equally confining for the ground and metastable states, the register traps are near a tune-out wavelength for the metastable states, providing near-zero trapping potential. Thus, any gate errors that lead to population being left in the metastable state (such as some coherent population errors on either stage of the two-photon transition) are converted to loss and can be detected (see Sec. A.6). This “erasure conversion” proves useful in identifying the location of errors WKPT (22).
We quantify the performance of two-qubit gates using several variations on randomized benchmarking. To isolate the performance of the gates themselves, without errors from atom movement, we perform interleaved randomized benchmarking (IRB) on static atoms within the IZ. In this protocol, we compare the return probability after a given number of random two-qubit Clifford operations (plus a return pulse calculated to return to a desired measurement state) to the same circuit with an additional CZ gate interleaved after each Clifford gate (and with an appropriately modified return pulse). By comparing the probability of returning to the measurement state versus the depth of the Clifford circuit with and without the added CZ gates, we isolate the error associated with the CZ gate alone, averaged over the two-qubit Hilbert space. Here, the atoms are kept in the IZ for all operations (including 1Q gates, unlike for typical circuits). Fits to IRB data taken out to depth 34 indicate an infidelity per CZ gate of 0.56(15)and additional loss and population leakage outside of the qubit subspace of 0.27(6)%.
In actual circuits, atoms are moved between the register and IZ between 1Q operations and 2Q operations. Because atom movement can cause heating, which may impact the performance of all following gates, IRB does not cleanly isolate the performance of the CZ gate when atoms are moved between arrays. For this, we rely on a variant of randomized benchmarking described in MSS+ (24) and similar to that in EBK+ (23), where concatenated circuit blocks each consist of a random 1Q operation and two CZ gates separated by an echo pulse. A final 1Q gate returns atoms to the measurement basis. We isolate the contribution from the 2Q gates (along with any loss or infidelity due to associated movements, up to single-qubit phases which are canceled by the echo pulses and characterized separately) by subtracting the per-depth infidelity associated with the 1Q operations. From this, we extract a per-CZ gate error of 0.39(10)and loss of 0.75(10)%. We perform this characterization with atoms originating in different rows within the register, as shown in Fig. 12, and do not observe a significant dependence on origin row or column. For comparison, the same procedure applied to atoms that remain static in the IZ for all operations yields a per-gate error rate of 0.35(5)and loss of 0.24(6)%.
Figure 12: Characterization of CZ gate performance on moved atoms. a, Echoed randomized benchmarking circuit applied with CZ gates omitted. Orange, green and blue points represent survival, fidelity for atoms that survive, and combined survival and fidelity, respectively. Lines represent exponential fits used to extract per-block error and loss rates. b, The same quantities with CZ gates added in (2 CZ gates per block). c, Measured infidelity (green) and loss (orange) for different qubit origin locations within the register. Statistical fluctuations may yield negative infidelity estimates because per-block error rates with CZ gates omitted are subtracted from those with CZ gates included.
A.6 Conversion of leakage to loss
Gate errors in the device, coherent or incoherent, can lead to population and phase errors within the qubit subspace, as well as leakage into other states. We operate the machine in a mode that converts population leakage from the qubit subspace into loss, which can then be identified independently from qubit state. This is particularly relevant in the context of 2Q gates, where population errors on the 1S0 to 3P0 or 3P0 to Rydberg state transitions can lead to population leakage. Conversion of these errors into loss is achieved by operating our register tweezers at a wavelength near 423 nm where 3P0 is nearly untrapped. (We could choose to depump the 3P0 states after each 2Q gate to avoid this loss, but find the conversion of leakage into a detectable error highly desirable.)
We characterize the effectiveness of the conversion from leakage to loss for 3P0 and the Rydberg state by preparing atoms in the register, moving them to the IZ, exciting them to the relevant state, moving them back to the register and then measuring the state and presence of the atoms. The metastable 3P2 states are anti-trapped in our IZ and rearrangement tweezers, and we do not repump these states. We observe greater than 99% of atoms to be lost when exciting to the clock state, and roughly 70% of atoms to be lost when exciting to the Rydberg state.
Appendix B Four-qubit codes
$\begin{array}[]{cccc}X&X&X&X\\ Z&Z&Z&Z\end{array}$
and logical operators
Observe that encoded ${|00\rangle}$ is the four-qubit cat state $\tfrac{1}{\sqrt{2}}({|0000\rangle}+{|1111\rangle})$. As shown in Fig. 5, an extra qubit is needed to prepare this state fault tolerantly Sho (96).
$S_{Z}XS_{Z}^{\dagger}$ $=Y$ $S_{Z}YS_{Z}^{\dagger}$ $=-X$ $S_{Z}ZS_{Z}^{\dagger}$ $=Z$
$S_{X}XS_{X}^{\dagger}$ $=X$ $S_{X}YS_{X}^{\dagger}$ $=Z$ $S_{X}ZS_{X}^{\dagger}$ $=-Y.$